OpenStudy (godlovesme):

HELP ASAP PLEASE!!!!! :(((((( Pyramid A is a square pyramid with a base side length of 18 inches and a height of 9 inches. Pyramid B has a volume of 3,136 cubic inches. How many times bigger is the volume of pyramid B than pyramid A? Give your answer as a percent. Provide an explanation and proof for your answer to receive full credit.

OpenStudy (godlovesme):

v= a^2 * h/3 or v= 1/3 * b * h im not sure which formula i have to use :/

OpenStudy (godlovesme):

@Nnesha help please :(

OpenStudy (godlovesme):

@LegendarySadist help please :))))

OpenStudy (anonymous):

Does it really matter which formula?

OpenStudy (godlovesme):

um you'll get different answer so i guess it does

OpenStudy (anonymous):

Not really. a^2 is the same as b. Since a is side length and b is base area. And area of a square is side squared. And 1/3 * h is the same as h/3. In this situation, it would be easier to use the first though imo.

OpenStudy (godlovesme):

if u use the first formula, you'll get 972 but if u use the 2nd one they you'll get 54 :/

OpenStudy (godlovesme):

then*

OpenStudy (anonymous):

Ok. The only difference is that the second formula is used when you have the BASE AREA and the first is used when you have the BASE SIDE LENGTH. The formulas themselves will get you the same thing, the only difference is the amount of steps. In the first formula, you'd just plug in 18 and 9 and calculate, while in the second, you'd find the base (18^2) and then multiply by 3. Both get the same result though.

OpenStudy (godlovesme):

ohh so which one do i use in this case? sorry im bad at this .-.

OpenStudy (anonymous):

You have the BASE SIDE LENGTH, so you'd use the first.

OpenStudy (godlovesme):

okay so it's gonna be 972 cuz v= 18^2 * 9/3 v= 324 * 3 v= 972

OpenStudy (anonymous):

Yes. The volume will be 972 in^3

OpenStudy (godlovesme):

yay! :) so what's the next step? :)

OpenStudy (anonymous):

\[\frac{3136}{972}\] Dividerino

OpenStudy (godlovesme):

3.226 since it's asking me for percentage, do i multiply it by 100?

OpenStudy (anonymous):

Yeah.

OpenStudy (godlovesme):

okay thank you soo much !! i really appreciate your help thanks! :)))

OpenStudy (anonymous):

No problem :)

OpenStudy (godlovesme):

(: